40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 10 03:19:58 2015, ericq, Update, LSC, AUX PDH error measured in CDS AUXspectra.png
    Reply  Fri Jun 12 17:09:58 2015, ericq, Update, LSC, Beatbox needs whitening gain beatGain.png
       Reply  Sat Jun 13 18:37:46 2015, ericq, Update, LSC, Beatbox needs whitening gain ADCdiagnosis.png
          Reply  Sat Jun 13 23:52:14 2015, ericq, Update, LSC, Beatbox needs whitening gain 
Message ID: 11351     Entry time: Wed Jun 10 03:19:58 2015     Reply to this: 11355
Author: ericq 
Type: Update 
Category: LSC 
Subject: AUX PDH error measured in CDS 

Looking over the old noise budget in the green locking paper, it seems the main technical noise sources were the AUX PDH error and DFD noise. I'm working on quantifying the current state of these noises. 

Rather than lugging out the analyzers every time I wanted to make a measurement of the AUX PDH error signals, I set out to make the existing digitized channels (ALS-[X/Y]_ERR_MON) usable for easier, and continuous, monitoring. Sadly, up until now the signals were poorly conditioned, and drowning in ADC noise. (When locked, the Y error signal was only +-10 counts!)

Of course, given the bandwidth of the green servos (10kHz), this won't tell us the full story of the what the green PDH lock is doing, but does indicate how much residual frequency noise exists in the ALS control band. 

I'm currently using SR560s at G=20 at each end to amplify the ERR MON outputs of the uPDH boards before sending them to the ADCs; now that I've found a gain that works, I'll modify the error point monitor buffer opamps inside the uPDH boards themselves during the daytime. 

The AUX Y error signal was going into an AA board with some funky filtering going on that I didn't want to mess with. Instead, I've moved the signal to the pentek generic board whose first four channels are used for the oplev segments, and the second quartet are unused, save for the TST channel I hooked up yesterday. 

On the pentek board, I changed the 4th order 800Hz lowpass to a 4th order 8kHz lowpass on the last three channels through some resistor swapping. (At first it was just the last two, but I found I was getting weird signals in the 32nd channel; and if I recall correctly from my cymac work, the 32 ADC channel is used for some timing signal or something...). I also turned off the 1:10 whitening filters on the last four channels via PCB jumper. 

I then unhooked the PZT drive and let the PDH error signals swing around, to calibrate the ADC counts into HZ. Now, the ALS-[X/Y]_ERR_MON_OUT_DQ are calibrated in green Hz! Here are the spectra.

As we've seen in the past (ELOG 10464), the X green is limited by the dark noise of the PD from 10-100Hz. This isn't so great. The RMS noise from 300Hz downwards (which is maybe the band where the ALS control would inject noise into the mirror motion) is about Y:10Hz X:40Hz.

During this time, the test masses were not under any longnitudinal control, so I'm not sure why there is such a difference in the height of the suspension resonance peaks, unless there's some differene in the low frequency PDH TFs that I've forgotten about. 

Now, with these references, we can easily check if the PDH loops change qualitatively over longer time periods.

I'll be including the effect of these noises in the upcoming revised ALS noise budget.

Attachment 1: AUXspectra.png  453 kB  | Hide | Hide all
AUXspectra.png
ELOG V3.1.3-